OpenStudy (anonymous):

Question number 2, @satellite73

OpenStudy (anonymous):

can't wait

OpenStudy (anonymous):

\[9x^2-16x+60=0\]\[ax^2+bx+c=0\] the distriminant is \[b^2-4ac\]which is our cases is \[(-16)^2-4\times 9\times 60\]

OpenStudy (anonymous):

if my arithmetic is right, this is \[(-16)^2-4\times 9\times 60=-1904\] does that look right?

OpenStudy (anonymous):

From what I know which is not much lol it looks pretty right to me (:

OpenStudy (anonymous):

k good ( i used a calculator) there are 3 possibilities for the discriminant a) is it positive meaning there will be two real solutions to the equation b) it is zero, which means there will be one real solution c) it is negative (our case) meaning there will be NO real solutions that is the answer h ere

OpenStudy (anonymous):

so when its negative there are no real solutions?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

how about \[4x^2+8x-5=0\] my suggestion is to factor it if you can

OpenStudy (anonymous):

Okay yeah i can do that but i have to explain as to why i chose to factor it could you tell me why you suggest it? To me it seems like the easiest method idk

OpenStudy (anonymous):

you want a hint? t is \[(2 x-1) (2 x+5) = 0\] factor because what you said, it is the easiest method if you can do it

OpenStudy (anonymous):

and also in this case you don't have too many choices because \(-5=-1\times 5\) or \(5\times (-1)\) so it is not hard to come up with \[(2 x-1) (2 x+5) = 0\]

OpenStudy (anonymous):

Okay thank you (:

OpenStudy (anonymous):

you can solve that, right?

OpenStudy (anonymous):

\[2x-1=0\\ 2x=1\\x=\frac{1}{2}\] for the first one

OpenStudy (anonymous):

yep i can (: and okay thanks

OpenStudy (anonymous):

good

OpenStudy (anonymous):

wait this is going to come out dumb i suppose but are we still on part b or did we go to C? sorry little tiny bit confused

OpenStudy (anonymous):

we did part B solved \[4x^2+8x-5=0\]

OpenStudy (anonymous):

\[(2 x-1) (2 x+5) = 0\] \[2x-1=0\\ 2x=1\\ x=\frac{1}{2}\] or \[2x+5=0\\ 2x=-5\\ x=-\frac{5}{2}\] are the two solutions

OpenStudy (anonymous):

okay yes got that, thanks for clearing up my confusion

OpenStudy (anonymous):

k now how about the last one i guess you can't factor it

OpenStudy (anonymous):

\[2x^2-12x+5=0\] we can complete the square

OpenStudy (anonymous):

subtract 5 from both sides to get \[2x^2-12x=-5\] then divide by 2 and ge t \[x^2-6x=\frac{5}{2}\] are the first two steps write them down, let me know when you want to continue

OpenStudy (anonymous):

okay done writing that down (:

OpenStudy (anonymous):

ok now we take half of 6, which is 3 and square it to get 9, and write \[(x-3)^2=-\frac{5}{2}+9\] or \[(x-3)^2=\frac{13}{2}\]

OpenStudy (anonymous):

maybe it would have been easier to use the quadratic formula, i don't know, but we are almost done

OpenStudy (anonymous):

okay i wrote both of them down

OpenStudy (anonymous):

you get \[x-3=\pm\frac{\sqrt{13}}{\sqrt2}\] so \[x=3\pm\frac{\sqrt{13}}{\sqrt2}\]

OpenStudy (anonymous):

next job i get will be doing math at mystic falls

OpenStudy (anonymous):

okay thankss (: and lol well that would be a little complicated cause mystic falls is actually a fictional town from the vampire diaires xD i actually go to FLVS

OpenStudy (anonymous):

yeah i know i googled it

OpenStudy (anonymous):

well, i didn't know about FLVS and don't even know what that is

OpenStudy (anonymous):

florida virtual school (: its home school but on a computer lol

OpenStudy (anonymous):

wow must be hard to learn math that way at least you can get help here do you have to turn this in? like written out on paper? or through a computer

OpenStudy (anonymous):

through the computer, we could mail our work to the teachers but it would take a while for it to get there so our teachers dont want us to.

OpenStudy (anonymous):

ok i assume you want to be done right? because we could redo the last one using the quadratic formula if you think it would look better to the teacher

OpenStudy (anonymous):

Oh no its fine, plus i kinda need to finish up cause its 11pm and i have to go to bed soon lol so can we go onto my last question?

OpenStudy (anonymous):

you got one more?

OpenStudy (anonymous):

cause that was the last one, the one where we got \[x=3\pm\frac{\sqrt{13}}{\sqrt2}\]

OpenStudy (anonymous):

oh yeah i know lol but i have one last question i need to work on if you cant help me now thats fine we can finish up tomorrow (:

OpenStudy (anonymous):

go ahead an post, maybe we can do it quickly